OpenStudy (meilendurcer):

1 3/4 divide 3= A. 20/4 B. 3 1/4 C. 1 4/7 D. 7/12 i have B, it's close the number i got.

zepdrix (zepdrix):

\[1\frac{ 3 }{ 4 }\] It might help to convert this mixed number to a fraction first =) \[\frac{ 4 }{ 4 }+\frac{ 3 }{ 4 }=\frac{ 7 }{ 4 }\] now our problem is \[\frac{ 7 }{ 4 }\div3\] At this point, I'm not sure what method you prefer to use to simply it, but you should get something like this :D \[\frac{ 7 }{ 4 }\div3=\frac{ 7 }{ 4 }\times\frac{ 1 }{ 3 }=\frac{ 7 }{ 12 }\]

OpenStudy (meilendurcer):

thank you so much!